OpenStudy (anonymous):

Hi! I have three questions about inequalities. I will give medal and fan to the best answer. Questions will be posted down below! thanks in advance

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } x \ge 12\]

OpenStudy (anonymous):

for the one that i just posted which statament goes with the equation One fourth times a number is greater than 12. One fourth times a number is less than 12. One fourth times a number is at least 12. One fourth times a number is at most 12.

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

if you divide both sides by 1/4 you will find your answer. Note on the right side, dividing by 1/4 is the same as multiplying by 4

OpenStudy (anonymous):

what? My answer doesnt match any of the options

OpenStudy (anonymous):

honestly, 1 PM is sufficient, if i'm on i will see it. if I am not helping someone else, I will come. Oh sorry, I solved for x. Answer me, how many x's do you have there?

OpenStudy (anonymous):

one

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

can u just quickly tell me the answer? I really need to turn this in!!

OpenStudy (anonymous):

a b c or d?

OpenStudy (anonymous):

I can't/won't do that. this question really... just look at the equation the answer is staring at you. We could have been done by now if you weren't resisting help.

OpenStudy (anonymous):

the answer is 

OpenStudy (anonymous):

ok I just have one question is this sign at most or at least?

OpenStudy (anonymous):

\[\le \]

OpenStudy (anonymous):

there is no math, it just asks you what the inequality is saying, read it

OpenStudy (anonymous):

You will need to know what greater than or equal signs mean that's all.

OpenStudy (anonymous):

One fourth times a number is at most 12.

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

that answer would be  so no